OpenStudy (anonymous):

Find the area of the region that is bounded by the given curve and lies in the specified sector

OpenStudy (anonymous):

\[r= 3+2 \sin \theta ,..... -\frac{ \pi }{ 2 }\le \theta \le \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

I will type what i have so far... give me a min plz

OpenStudy (anonymous):

\[A= \int\limits_{\frac{ -\pi }{ 2 }}^{\frac{ \pi }{ 2 }}\frac{ 1 }{ 2 }(3+2 \sin \theta)^2 d \theta\]=\[A= \frac{ 1 }{ 2 }\int\limits_{\frac{ -\pi }{ 2 }}^{\frac{ \pi }{ 2 }}(9+ 12 \sin \theta + 4 \sin^2 \theta) d \theta\]=\[A= \frac{ 1 }{ 2 }\int\limits_{\frac{ -\pi }{ 2 }}^{\frac{ \pi }{ 2 }}(9 +4 \sin^2 \theta) d \theta\]=\[A= \frac{ 1 }{ 2 }(2)\int\limits_{0}^{\frac{ \pi }{ 2 }}( 9 +4 \sin^2 \theta) d \theta\]

OpenStudy (anonymous):

I know a half angle identity goes in at this point, but this is where i get messed up I also know that the answer is \[\frac{ 22\pi }{ 4 }\]

OpenStudy (anonymous):

the sin^2 theta needs to becom \[\frac{ 1 }{ 2 }(1-\cos 2 \theta )\] and this messes me up

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{ \pi }{ 2 }}( 9 +4 [\frac{ 1 }{ 2 }(1-\cos 2 \theta)] )d \theta\]

OpenStudy (tkhunny):

Not sure I would do that: \(\int \left(3+2\sin(\theta)\right)^{2}\;d\theta = \theta\cdot\left(3+2\sin(\theta)\right)^{2} - \int \theta\cdot 2\cdot\left(3+2\sin(\theta)\right)\cdot\cos(\theta)\;d\theta\) After that, your half angle looks a lot more obvious.

OpenStudy (anonymous):

I dont know how to do your way, im sorry

OpenStudy (freckles):

he did integration by parts

OpenStudy (anonymous):

oh...

OpenStudy (freckles):

which looks really cute too

OpenStudy (anonymous):

that explains a lot... i suck at that the most

OpenStudy (freckles):

though I think he is missing a multiple of 2 in the integral part

OpenStudy (anonymous):

btw, hi freckles! its been a min

OpenStudy (anonymous):

I would prefer to continue the way i started, its easier for me

OpenStudy (freckles):

and hey back at ya

OpenStudy (freckles):

that's fine too

OpenStudy (anonymous):

but then again, I am stuck at where i posted... the algebra part

OpenStudy (freckles):

this is what you posted \[\int\limits_{0}^{\frac{ \pi }{ 2 }}( 9 +4 [\frac{ 1 }{ 2 }(1-\cos 2 \theta)] )d \theta\] i think doing a little distribution would make this looks tons prettier

OpenStudy (freckles):

first of all 4/2 =2 so you actually have for that mumbo jumbo part the following 9+2(1-cos(2theta)

OpenStudy (freckles):

I forgot to close that

OpenStudy (anonymous):

I thought that too,.. but where does the 4 come from in the answer?

OpenStudy (freckles):

but you can now distribute that two there

OpenStudy (freckles):

oh are you following an example from the book? this isn't your work?

OpenStudy (anonymous):

this is my homework online, it gives me the answer...

OpenStudy (anonymous):

well, i chose the correct multiple choice at least... the answer is 22pi/4

OpenStudy (freckles):

like as it shows you the work and you are trying to follow it?

OpenStudy (anonymous):

it doesnt show me the work, i just chose the right answer :(

OpenStudy (freckles):

lol sorry but you came up with this \[\int\limits_{0}^{\frac{ \pi }{ 2 }}( 9 +4 [\frac{ 1 }{ 2 }(1-\cos 2 \theta)] )d \theta\] and you are asking me where you got the 4 from?

OpenStudy (anonymous):

no, the answer is \[\frac{ 22\pi }{ 4 }\] i was wondering about the 4 in the denom

OpenStudy (anonymous):

i know im close here, i just get stuck when integrating here...

OpenStudy (freckles):

\[\int\limits_{0}^{\frac{\pi}{2}}(9+\frac{4}{2}(1-\cos(2 \theta)) d \theta \\ \int\limits_0^\frac{\pi}{2}(9+2 (1-\cos(2 \theta)) d \theta \\ \int\limits_0^\frac{\pi}{2}(9+2 -2 \cos(2 \theta)) d \theta \\ \int\limits_0^\frac{\pi}{2}(11-2 \cos(2 \theta)) d \theta \] so you got to this part right ?

OpenStudy (anonymous):

oooh... nope, i didnt do that far, i got confused there

OpenStudy (freckles):

let me know if you don't understand something I did there with the distribution part

OpenStudy (freckles):

if you have no questions with that part can I see your antiderivative

OpenStudy (anonymous):

I see that now! but i forget how to do the integration, is it \[11 \theta - 2(\frac{ 1 }{ 2 }) \cos 2 \theta \]

OpenStudy (freckles):

almost use the the antiderivative of cos(x) is sin(x)

OpenStudy (anonymous):

oh yea...

OpenStudy (anonymous):

oh yea...\[A= 11 \theta - \sin 2 \theta \rightarrow \frac{ \pi }{ 2 }\]

OpenStudy (freckles):

right \[[11 \theta-\cancel{2} (\frac{1}{\cancel{2}} ) \sin(2 \theta)]_0^\frac{\pi}{2}\]

OpenStudy (freckles):

and yeah you only need to plug in the upper because the lower makes it all 0

OpenStudy (anonymous):

ok so...

OpenStudy (anonymous):

\[11(\frac{ \pi }{ 2 })- \sin (\frac{ \pi }{ 2 })= \frac{ 11\pi }{ 2 }-1\]

OpenStudy (freckles):

well the only problem with that is you had a 2 inside that sin thingy

OpenStudy (freckles):

2*pi/2=pi

OpenStudy (freckles):

sin(pi)=0

OpenStudy (freckles):

so you have 11pi/2

OpenStudy (freckles):

multiply top and bottom by 2 and you have your desired unsimplified answer

OpenStudy (anonymous):

ok... where does the 22pi/4 come from then?

OpenStudy (anonymous):

oh... i guess i dont see where the multiplying by 2 top and bottom came from

OpenStudy (freckles):

well 2/2=1

OpenStudy (anonymous):

yes.. OH , where we simplified above?

OpenStudy (freckles):

\[\frac{11\pi}{2} \cdot \frac{2}{2}=\frac{22 \pi}{4}\] this answer they have is simplified 11pi/2 is actually the simplified version of the answer don't ask me why they prefer the other :p

OpenStudy (anonymous):

I bet you if this wasnt multiple choice and i plugged in 11pi/22, it would say correct!

OpenStudy (freckles):

this answer they have is un-simplified*

OpenStudy (anonymous):

weird

OpenStudy (freckles):

you mean 11pi/2

OpenStudy (anonymous):

lol thanks AGAIN

OpenStudy (freckles):

np

OpenStudy (anonymous):

I may have another.... well c

OpenStudy (freckles):

ok i might be around a little while longer

OpenStudy (tkhunny):

Post on another thread. Get better at Integration by Parts. :-)